How many moles of Cl- ions are present in 259 mL of a 0.16 M FeCl3 solution


1
Expert's answer
2021-03-09T06:04:02-0500

The formula tells you that there are 3 mol of Cl⁻ in 1 mol of FeCl3. So, we must find the moles of FeCl3.

Proportion:

0.16 mol – 1000 mL

x mol – 259 mL

"x = \frac{0.16 \times 259}{1000}= 0.0414 \;mol"

n(FeCl3) = 0.0414 mol

n(Cl⁻) "= 3 \times 0.0414 = 0.1242 \;mol"

Answer: 0.1242 mol